OpenStudy (justin528p):

@tom982 help 1.Graph the following conditions. Click on the graphic to choose the correct answer. {(x, y): x + y = 5} 2.Indicate in standard form the equation of the line passing through the given point and having the given slope. R(4, 0), m = 5

OpenStudy (justin528p):

its a test cant fail this is my final test for the semester i cant retake this class D:

OpenStudy (justin528p):

@tom982

OpenStudy (justin528p):

these r diff the first one involves a graph

OpenStudy (anonymous):

What do you think the answer is? What is the gradient of that line? In other words, how much is y going up by every time x increases by 1.

OpenStudy (justin528p):

y goes up 1

OpenStudy (justin528p):

@tom982 ??

OpenStudy (anonymous):

Yes, so we know m=1. What is the y intercept?

OpenStudy (justin528p):

5 :D

OpenStudy (justin528p):

x is -5 so what do i do to set it up @tom982

OpenStudy (justin528p):

?? any1 :( i got 20 mins

OpenStudy (anonymous):

5 is correct. So our line is y=x+5. Therefore the graph you sent me doesn't show {(x, y): x + y = 5}. You're looking for the graph of x+y=5, or y=-x+5, which looks like this: 

OpenStudy (justin528p):

this is the graph they gave me https://tiftschoolsga.owschools.com/media/g_geo_2015/9/qu3025c.gif

OpenStudy (justin528p):

oh nvm i got it thx again @tom982 !!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

No problem, happy to help.

OpenStudy (justin528p):

:D